DATES PRICE INDEX
1974-01-01 26.574
1974-02-01 26.883
1974-03-01 27.195
1974-04-01 27.387
1974-05-01 27.657
1974-06-01 27.891
1974-07-01 28.098
1974-08-01 28.391
1974-09-01 28.687
1974-10-01 28.893
1974-11-01 29.110
1974-12-01 29.341
1975-01-01 29.523
1975-02-01 29.684
1975-03-01 29.786
1975-04-01 29.887
1975-05-01 30.007
1975-06-01 30.195
1975-07-01 30.454
1975-08-01 30.586
1975-09-01 30.737
1975-10-01 30.916
1975-11-01 31.114
1975-12-01 31.286

a) (5 points) Calculate the rate of inflation from January 1974 (1974-01-01) to January 1975 (1975-01-01).

b)(5 points) Suppose that you earned $30,000 per year back in January of 1975. Assuming that the current price index is 109, how much money would you have to make now to have the same purchasing power as $30,000 did back in January of 1975?

c) (5 points) In January 1974, the one year interest rate 7.01%. Calculate the ex-post real rate of interest.

d) (5 points) Given your answer in part c), is this an environment attractive to saving or spending? Explain.
Answer #1

A) Inflation from 1974-1975 = (29.523 - 26.574)/26.574 = 0.1109 or 11.09%

B) Index in Jan 1975 = 29.523

Income in 1975 = 30,000

Index today = 109

Thus, income today = (30,000/29.523) x 109 = $110,761.1

C) Inflation in Jan 1975 = 11.09%

Nominal interest rate = 7.01%

Ex post real interest rate = 7.01 - 11.09 = -4.08%

D) This is an attractive time to spend, as real interest rate is negative